name2004-iss3: Logbook Entries

name2004-iss3: INFO Messages: 13 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
191 Wed 11-Aug-2004P. Penasco Motosat problemsISS2 (Puerto Penasco)Lou
190 Wed 11-Aug-2004Kino Roving team visitISS3 (Bahia Kino)June
145 Mon 02-Aug-2004Clary UPS malfunctionRebeca
132 Thu 29-Jul-2004Roving team COMMENTS ON LOS MOCHISISS3 (Bahia Kino)Steve/John
131 Thu 29-Jul-2004Roving team notes for July 29ISS3 (Bahia Kino)Steve/John
120 Tue 27-Jul-2004Roving team notes from LORETO visitISS3 (Bahia Kino)Steve/John
119 Tue 27-Jul-2004Roving team notesISS3 (Bahia Kino)Steve/John
99 Thu 22-Jul-2004Roving team notes: Thursday July 22ISS3 (Bahia Kino)Steve/John
78 Sun 18-Jul-2004Roving team report for July 18ISS3 (Bahia Kino)Steve/John
49 Mon 12-Jul-2004Roving team notesISS3 (Bahia Kino)Steve/John
12 Mon 05-Jul-2004training sessionISS3 (Bahia Kino)Operator
10 Sun 04-Jul-2004training sessionISS3 (Bahia Kino)Charlie
3 Tue 29-Jun-2004Its AliveISS3 (Bahia Kino)Lou

191: INFO, Site ISS2 (Puerto Penasco), Wed 11-Aug-2004 19:07:47 GMT, P. Penasco Motosat problems
Since yesterday afternoon I've been working with the operators at Penasco to fix the motosat system. In theory the system should be working. All tests have have worked fine . System passed calibrate sensors ,good sat numbers for strengh and quality. System has been shutdown and rebooted ,dish stowed and brought up
 again. GPS up dates when refreshed.
Still time out error messages when attempting to go to any web site.
 Currently waiting on a return call from Eagle Computers to discuss what could 
be happening.
 From what operators have said porblems started after some disruption in power
to the system.

190: INFO, Site ISS3 (Bahia Kino), Wed 11-Aug-2004 18:51:16 GMT, Kino Roving team visit
The roving team (Lou and June) visited Kino site on Aug. 10 and Aug. 11. Here is a summary for our visit:

(1) Profiler data backup: We stopped the profiler at 20:05-20:55 UTC on Aug. 10 to back up the profiler data. The data were backed up to two CDs (NAME_KIN04 and NAME_KIN05). We left one copy in the trailer and brought one copy with us. After backup, the disk was ~20% full.

(2) Data backup on DM computer: We backed up raw and NIMA data on DM computer on two DVDs (one for raw and one for NIMA), left one copy in the trailer and brought one copy with us.

145: INFO, Site , Mon 02-Aug-2004 10:51:31 GMT, Clary UPS malfunction
When we tried to turn on the clary ups this morning, we realized it was not working so we  removed the plug and plugged it into the protection box as told by 
the roving crew on their last visit.
132: INFO, Site ISS3 (Bahia Kino), Thu 29-Jul-2004 20:51:48 GMT, Roving team COMMENTS ON LOS MOCHIS
Some good some bad:

 - Brad reported that the Motosat was set to "Do Not Disturb" which is why we
   could not call in from the other sites.
 - Brad also found the Fluke digital multimeter which the roving crew was
   falsely accused of removing.

 - Two 408 sondes launched at 04 and 08 Z on July 29 didn't look good. There
are not data from ~700-180 mb for 08Z and from ~250-130 mb for 04 Z. But it
was not documented in the tklog file. The students need to pay attention!

 - At least some of the Los Mochis crew have been trained on entering
   coefficients by hand from the paper tape.

131: INFO, Site ISS3 (Bahia Kino), Thu 29-Jul-2004 20:36:44 GMT, Roving team notes for July 29
Today is the last day of roving duty for Steve and John. Lou and June are
here in Kino now and the plebian order of the cell phone will be bestowed
upon them in a ceremony at dinner.

IOP#4 will finish this evening.

Tasks today:

- We did another data backup to DVD of the ISS-3 workstation.
- Showed the new roving team everything we could think of

Things are looking up at Los Mochis. Brad reported that the Motosat was set to
"Do Not Disturb" which is why we could not call in from the other sites. Brad
also found the Fluke digital multimeter which the roving crew was falsely
accused of removing.
At least one or two of the 408 sondes has launched fine.
At least some of the Los Mochis crew have been trained on entering coefficients
by hand from the paper tape.

The GLASS system has been put back on the CLARY GT1000 ups which isw on 3 extension cords .
The operators have been told that if there are any problems to remove plug strip
and to put the strip into the protection box.This is because of high voltage. 
The fix sent to us by CLARY is for different model. The 1000's input voltage 
is not adjustable.

to be continued....

120: INFO, Site ISS3 (Bahia Kino), Tue 27-Jul-2004 20:36:46 GMT, Roving team notes from LORETO visit
The LORETO Glass station does not use TKLOG so we are entering a short summary
from here at ISS-3/Kino.

Roving team (Steve and John) flew Hermosillo to Loreto last Friday (July 23).
The Glass site is at the Hotel Pinta which is booked for the season.
Hotel Posada de Loreto was fine. Basic, clean, 395 pesos. It is in the
center of town and a 10-minute walk from the Pinta.

A car is not needed except for transportation to/from the airport. Katie
gave us a ride, but a taxi would have worked.

Flight info:
Aeromexico has one flight/day same each day:
 HMO to LOR: Flight AM2290 departs at 4:35pm
 LOR to HMO: Flight AM2291 departs at 9:10am
 Price was about 3700 pesos round trip (depends on day of week)


 The Motosat was not working when we arrived. It had failed following a quick
manual shutdown during a strong windstorm. After many restarts by the crew over
many days it got to the configuration screen. Per conversation with Tim we used
the point4020 software on another PC (Steve's laptop on the LAN) to set the
satellite lat/lon. It has worked ok since. A continuing issue is that there are
no other PC's on that LAN which can successfully run point4020.

 The high-pressure gauge on the Loreto regulator was inoperative, and the hose
may leak. We brought the spare regulator from Puerto Penasco. The low pressure
hose was moved from the Loreto regulator. This regulator will now screwed on
the tanks without the hose. Appears to work well, and without leaking. We also
played with the balloon size. In Kino the students measured a balloon
circumference with 34 cf. The students at Loreto cut a rope to this length.
When filled with 400 lbs helium, the curcumference matched that in Kino

 The Glass PC failed after the 12Z sounding on July 25. It appeared to be
a hard drive failure. The students told us that the setup crew had much
trouble with this PC. With reckless abandon JWM opened the laptop and noted
the HDD was loose in its tray. Further there were no screws holding
the HDD in the tray and only 1 screw holding the tray to the case.
A cardboard shim was fashoned
to reduce play of the HDD in its bracket. THe HDD was reinstalled with
a satisfying connection to the motherboard. The PC has been better behaived
since. We ran scandisk which found many "errors". Some were related to old
software (versions of the GLASS software and others). After consultation with
Tim, and over two days, we removed considerable software from this PC. We then
began to defragment the disk. This process had to be interupted several times
to launch sondes. The crew will continue to defrag between launches until
it is done. Apologies to the crew in Boulder for calling an emergency and then
fixing the problem. Ned was willing to scramble an F-18 to get us a replacement

A second problem with this PC was noted: The case fan is not working.
We are now using four Glass pylons to lift the Glass pc above the desk,
providing the possibility of ventilation below. We have also bought a fan for
the room to circulate air past the PC.

Lou is bringing a replacement pc for GLASS and should install it soon. We would
have tried running GLASS on another laptop but apparently the software can
only be run under Windows-98.

 The same storm that caused the Motosat to be stowed blew down the GLASS
surface met tower. The students righted the tower and replaced it in the same
position. We tied the station to two nearby trees (in addition to the stakes
already in the sand at the tower base). It turns out that this tower had not
been working since setup. There was no communication between the GLASS pc and
the campbell. JWM worked on that and eventually made contact. But the values
coming back were clearly not calibrated. Possibly we were using the wrong DLD
file? Because we were concerned with data quality/accuracy and did not think
a continuous data stream would benefit the PI we did not pursue this further.
Data quality issues:
 - the anemometer propellor broke when the station was blown over
 - the pressure sensor has no static port, so P is affected by wind around the
Campbell box.
 - the T/RH sensors are aspirated but the fan is weak.
 - numerous trees and large buildings were clearly affecting the local flows
at the tower height.
 - the anemometer orientation was off >=8 degrees and the reported direction
was even further off indicating that the potentiometer may be rotated in the
collar, or the 'BD or ZN' boom direction/zero north are incorrect.

 GLASS expects to use the surface tower for a ground point (P,T,H,WSpd,Wdir),
not for regular time series data. Since this was not available they are simply
copying the sonde values for surface data, estimating the wind speed, and
using the anemometer vane to estimate direction. (Note that we measured the
tower black box direction to be about 8 degrees clockwise of true south).

 Be aware that there is NO test equipment, few tools, and no manuals at this


-- We noticed that GLASS does not aspirate the rawinsonde T/RH sensor prior to
launch. The ISS group has found this to be necessary, especially if the sensor
arm is in direct sunlight.

-- The GLASS receiving antenna is poorly positioned at this site. Its view of
the sky is blocked to the west to more than 45 degrees above the horizon by
the hotel building. This is because the coax between the GLASS rack and GLASS
amplifier, and between the amplifier and the antenna are fairly short. We saw
flights where the balloon went west and we lost signal around 300 mb because
the sonde signal was blocked by the building. We've recommended the next Roving
team bring a replacement/longer cable. Lou reported that a longer coax had to
be used for the GPS ground antenna in this installation.

119: INFO, Site ISS3 (Bahia Kino), Tue 27-Jul-2004 20:30:46 GMT, Roving team notes
We (Steve and John) are back in Kino.

IOP#4 is due to begin this afternoon.

Did a profiler backup of days 200-208 onto CD NAME_KIN03.
The profiler was down from 1921-2022Z.

Did a data backup DVD of the ISS-3 workstation. Will bring back to Boulder.


Hotel: we are trying the 

99: INFO, Site ISS3 (Bahia Kino), Thu 22-Jul-2004 18:34:47 GMT, Roving team notes: Thursday July 22
Steve and John are stopping by to drop our luggage before a flight to Loreto
this afternoon.

The profiler failed sometime around 17:30. It appeared the TX was not
transmitting. RX sample comb was ok. Solved by powering down RX/TX and
interface boxes, rebooting PC. Checked cable connections. Then it worked.
Good news since our flight leaves in a couple of hours.

All else looks great here.

78: INFO, Site ISS3 (Bahia Kino), Sun 18-Jul-2004 19:40:09 GMT, Roving team report for July 18
We are back in Kino for a day.

Backed up Profiler SPC files onto two CD's. Two copies of each, one for
the trailer and one to take home. The profiler was down for an hour during
the process. 1705-1803Z. Note that there is about 1.5GB free so good for
almost another 3 weeks. Disk is about 19% full (there is a mystery of what
is taking up about 400 MB of space on D:)

Checked the voltage on the CLASS pre-amp under load. 11.37 volts which is
fine. Much better than Los Moches was. ALso checked T-dome and T-case wiring
is correct. LATER: Checked the pre-amp again during a sounding and found
a lower voltage. Some effect of having the RF in the circuit? Conclusion is
is not to change the class receiver.

Line voltage to ISS is still too high at 134.2 V. JWM set up a workaround.
We have the UPS preceeded by two 100-foot extension cords to drop the voltage
to within the UPS limits.

Ran data-backup script again. Taking DVD back to Boulder.

49: INFO, Site ISS3 (Bahia Kino), Mon 12-Jul-2004 19:38:32 GMT, Roving team notes
Steve, John, Charlie at Kino today. Hal flew home this morning.

We got to meet Jennifer and Brad. Sounding from here look very good.
They note a couple of times losing signal when the sonde travels far
to the SSE. No big obstructions in thtat direction.

We had an impressive lightning show at the Hermosillo hotel last
night, from about 9-9:45pm (4-4:45Z). The storm was moving East to West.

The Bahia Kino soap plots show rain aloft approx. 5:45-7:00Z and another
brief shower around 8:30Z.

Rain gauge shows rain at the surface: 0.4 mm near 6Z and another 0.1 mm near 7Z.
No tips around 8:30Z.

Profiler POP consensus winds look good, despite some bird clutter in the
individual dwells. Nice match of low winds with surface tower.

We added tie-downs to the rain gauge.
Checked profiler level: ok. In fact about as level as can be.

12: INFO, Site ISS3 (Bahia Kino), Mon 05-Jul-2004 00:45:38 GMT, training session
Conducting Bahia Kino training session. Initiated Power Off at approx 530 LST, repowered system at 545 LST...everything fine...

10: INFO, Site ISS3 (Bahia Kino), Sun 04-Jul-2004 21:49:02 GMT, training session
Conducting Bahia Kino training session.

3: INFO, Site ISS3 (Bahia Kino), Tue 29-Jun-2004 23:56:20 GMT, Its Alive
ISS3 Kino Bay Is alive